A ball of mass 8kg falls from rest of a height of 100m. Neglecting air resistance, calculate it's kinetic energy after falling a distance of 30m.(Take g as 10ms^-2).


1
Expert's answer
2021-02-17T10:41:42-0500
"KE=\Delta PE=mg(h_2-h_1),""KE=8\ kg\cdot10\ \dfrac{m}{s^2}\cdot(100\ m-70\ m)=2400\ J."